# Prompt to Restore, Restore to Prompt: Cyclic Prompting for Universal Adverse Weather Removal

### v1

@@ IV-C Comparison with State-of-the-art Methods
TABLE II: Quantitative Results under Real-World Scenarios. Text Denotes the Best Performance and Underline Denotes the Second Best. The FLOPs are computed on the input image of size $256\times 256$.
Methods SPA [88] RainDS [28] Snow100K-real [9] Params. GFlops
@@
TransWeather [15] 27.96 0.8677 45.83 21.78 0.690 49.64 60.56 0.388 3.042 37.7 6.13
WeatherDiff128 [27] 29.26 0.9279 46.32 22.40 0.6405 59.00 62.48 0.396 2.970 85.6 65.9$\times$169$\times$25
T3-DiffWeather [26] 31.36 0.9340 45.01 23.12 0.6461 45.01 61.65 0.397 2.910 69.4 59.8$\times$2
T3-DiffWeather (4 steps) [26] 31.65 0.9367 46.10 23.17 0.6388 43.58 61.70 0.397 2.922 69.4 59.8$\times$4
GridFormer [87] 31.18 0.9269 46.23 22.89 0.6377 46.85 61.49 0.395 2.912 30.1 251.35
CyclicPrompt (ours) 31.95 0.9347 46.51 23.10 0.6620 50.32 61.72 0.395 2.893 29.9 230.42
@@
Quantitative Comparison. Table I provides the quantitative results of all methods across three adverse weather types. We can see that the task-specific methods, e.g., MPRNet [85] and IDT [46], produce promising results on the weather type when applying their corresponding customized models or even better than several UAWR methods (e.g., All-in-one [14]). However, as shown in Table I, they fail to be generalized to other out-of-distribution weather types, such as MPRNet on raindrops and IDT on snow, where their PSNR and SSIM decrease significantly. When we retrain the general restoration methods on the same All-Weather dataset according to the UAWR setting, one can observe that the resultant models show better generalizability on three different datasets (MPRNet v.s. MPRNet*). Our CyclicPrompt outperforms existing methods by a significant margin across almost all weather types and achieves the best on average. Compared to the recent most SotA method T3-DiffWeather, our method exceeds it over 0.3dB in average PSNR. This is mainly due to our holistic prompt construction and distinctive cyclic prompt learning design, which is preferable for UAWR.
@@ Methods
Qualitative Comparison. In Figure 4, we visualize the restoration results of all compared methods. It can be observed that, under various weather conditions, our method can consistently remove the adverse effects of unfavorable weather from corrupted images and preserve plausible textures. Especially, in the rain+fog degradation, our method effectively recovers obscured scenes with more fine-grained details.
Real Scenario Comparison. We also evaluate the performance of our CyclicPrompt in real-world scenarios under different weather conditions and compare it with more recent methods. We use the SPA [88], RainDS [28], and Snow100K-real [9] datasets as the testing sets. We measure the performance using PSNR, SSIM, and a non-reference metric MUSIQ [105]. For Snow100K-real, since the ground truth images are unavailable, we apply the non-reference metrics MUSIQ, MANIIQA [106], and NIQE [107] for quantitative evaluation. Besides, the model complexity measured by parameters (Params., M) and computational costs (GFLOPs, G) is also reported. As illustrated in Table II, Transweather requires the fewest computations but more parameters than CyclicPrompt and suffers from significant performance reduction. T3-DiffWeather performs efficiently but with over 2×2$\times$2 × parameters and slightly lower scores than CyclicPrompt. To make a more intuitive comparison, we increase the iteration step of T3-DiffWeather to 4, ensuring it has similar computational costs (239.2G) to CyclicPrompt. The results reveal that the model does not yield apparent improvements, which still underperforms CyclicPrompt across most metrics. GridFormer shows relatively heavier model complexity but is not adept in real-world weather, which performs worse. Though WeatherDiff128 is comparable to ours, this method involves about 2.5×2.5$\times$2.5 × more parameters and 1200×1200$\times$1200 × more FLOPs. CyclicPrompt achieves competitive performance with obviously smaller model complexity.
@@ IV-D Ablation Study
Effect of C2P. As discussed in Sec. III-B, the initial C2P consists of weather-specific knowledge, learnable input-conditional vector, and textual prompt. We denote the variant without any prompt as the baseline model. As shown in Table IV, to verify the effectiveness of the initial C2P, we first induce the weather-specific knowledge in the baseline, which exploits the pre-trained CLIP image encoder to extract visual degradation features, prompting the model to improve. The input-conditional vector can be seen as a dynamic prompt for input content, which can calibrate and adapt the weather-dedicated information to complex weather, thereby providing a favorable gain of 0.21dB. Additionally, in Fig. 6, we conduct visual comparisons between the models with and without the input-conditional vectors, where the results validate that the latter can better remove the weather. Interestingly, the textual prompt corresponding to the weather scene is beneficial for restoration as it can incorporate explicit scene contexts and semantics, even if it stems from low-quality observations. Hence, the model with full C2P achieves the best.
@@
TABLE V: Ablation Study on the Number of Input-conditional Vector.
Vector Number
0 4 8 12
PSNR 30.83 30.87 31.18 30.37
SSIM 0.9217 0.9227 0.9251 0.9222
@@
Number of Learnable Vectors. The input-conditional vector in C2P is of the size $N\times C$. To further study the impact of the learnable vectors for adverse weather removal, we initialize the vectors with different numbers (i.e., $N$), where “0” denotes the model without the vector. The results are summarized in Table V. We can see that the performance boosts gradually as the number increases. This is due to more vectors yielding better capacity for context understanding and weather-related prompt adaption. However, when we increase $N$ from 8 to 12, the model suffers from performance decreases. This can be attributed to the overlarge vectors introducing unnecessary noisy information that overwhelms other important components. Therefore, we set $N=8$ in our experiments.
@@
In CyclicPrompt, the residual prior is incorporated to enrich visual characteristics for better detail recovery along with EPM. We now present the experimental results to verify its impact. First, in Table VI, we observe that it assists well the model and enhances the performance notably. Moreover, in Fig. 7, we visualize the residual prior map and corresponding feature maps before and after the RPM. There are two observations: (1) the prior (Fig. 7(b)) is weather-invariant and allows expressing the contextualized information of scenes. (2) By modulating the decoder feature (Fig. 7(c)) with the prior, the model can effectively emphasize the textures and structures (Fig. 7(d)), improving the representation capacity for reconstruction.

### v2

@@ IV-C Comparison with State-of-the-art Methods
TABLE II: Quantitative Results under Real-World Scenarios. Text Denotes the Best Performance and Underline Denotes the Second Best. The FLOPs are computed on the input image of size $256\times 256$. We calculate the average FPS on the RainDS ($1296\times 768$) and SPA ($512\times 512$) datasets.
Methods SPA [89] RainDS [28] Snow100K-real [9] Params. GFlops FPS
@@
TransWeather [15] 27.96 0.8677 45.83 21.78 0.690 49.64 60.56 0.388 3.042 37.7 6.13 56.7784
[…]
GridFormer [88] 31.18 0.9269 46.23 22.89 0.6377 46.85 61.49 0.395 2.912 30.1 251.35 2.1546
CyclicPrompt (ours) 31.95 0.9347 46.51 23.10 0.6620 50.32 61.72 0.395 2.893 29.9 230.42 2.5231
Quantitative Comparison. Table I provides the quantitative results of all methods across three adverse weather types. We can see that the task-specific methods, e.g., MPRNet [86] and IDT [47], produce promising results on the weather type when applying their corresponding customized models or even better than several UAWR methods (e.g., All-in-one [14]). However, as shown in Table I, they fail to be generalized to other out-of-distribution weather types, such as MPRNet on raindrops and IDT on snow, where their PSNR and SSIM decrease significantly. When we retrain the general restoration methods on the same All-Weather dataset according to the UAWR setting, one can observe that the resultant models show better generalizability on three different datasets (MPRNet v.s. MPRNet*). Our CyclicPrompt outperforms existing methods by a significant margin across almost all weather types and achieves the best on average. Compared to the recent most SotA method T3-DiffWeather, our method exceeds it by over 0.3dB in average PSNR. This is mainly due to our holistic prompt construction and distinctive cyclic prompt learning design, which is preferable for UAWR.
Besides, we also implement two model variants by: 1) removing the textual prompt (w/o text), thus the model relies solely on visual prompts and learnable conditional vectors; 2) randomly shuffling the paired textual descriptions in the test set (w/ random text), feeding non-corresponding captions as textual prompts to the model. In Table I, we can see that removing textual prompts causes only a slight performance drop, indicating that the model effectively leverages visual and learnable prompts to maintain restoration quality. Though mismatched textual prompts lead to moderate degradation, it still performs competitively. Moreover, all our model variants achieve comparable performance across most datasets and superior performance on average against existing methods, demonstrating the robustness of our method in practical applications.
Qualitative Comparison. In Fig. 5, we visualize the restoration results of all compared methods. It can be observed that, under various weather conditions, our method can consistently remove the adverse effects of unfavorable weather from corrupted images and preserve plausible textures. Especially, in the rain+fog degradation, our method effectively recovers obscured scenes with more fine-grained details.
@@ Methods
Real Scenario Comparison. We also evaluate the performance of our CyclicPrompt in real-world scenarios under different weather conditions and compare it with more recent methods. We use the SPA [89], RainDS [28], and Snow100K-real [9] datasets as the testing sets. We measure the performance using PSNR, SSIM, and a non-reference metric MUSIQ [106]. For Snow100K-real, since the ground truth images are unavailable, we apply the non-reference metrics MUSIQ, MANIIQA [107], and NIQE [108] for quantitative evaluation. We also measure the model complexity by calculating the model parameters (Params., M), computation costs (GFLOPs, G), and average inference speed in FPS on two datasets: RainDS ($1296\times 768$) and SPA dataset ($512\times 512$).
As illustrated in Table II, Transweather requires the fewest computations with the fastest speed but more parameters than CyclicPrompt and suffers from significant performance reduction. T3-DiffWeather performs efficiently but with over $2\times$ parameters and slightly lower scores than CyclicPrompt. To make a more intuitive comparison, we increase the iteration step of T3-DiffWeather to 4, ensuring it has similar computational costs (239.2G) to CyclicPrompt. The results reveal that the model does not yield apparent improvements, which still underperforms CyclicPrompt across most metrics. We can also see that its efficiency is slowed down, which shows fewer FPS than ours. GridFormer shows relatively heavier model complexity but is not adept in real-world weather, which performs worse and more slowly than our method. Though WeatherDiff128 is comparable to ours, this method involves about $2.5\times$ more parameters and $1200\times$ more FLOPs. CyclicPrompt achieves competitive performance with obviously smaller model complexity and over $200\times$ FPS.
@@ IV-D Ablation Study
Effect of C2P. As discussed in Sec. III-B, the initial C2P consists of weather-specific knowledge, learnable input-conditional vector, and textual prompt. We denote the variant without any prompt as the baseline model. As shown in Table IV, to verify the effectiveness of the initial C2P, we first induce the weather-specific knowledge in the baseline, which exploits the pre-trained CLIP image encoder to extract visual degradation features, prompting the model to improve. The input-conditional vector can be seen as a dynamic prompt for input content, which can calibrate and adapt the weather-dedicated information to complex weather, thereby providing a favorable gain of 0.21dB. It can also be seen that the two components cooperate well and achieve about 0.3dB gains in PSNR, which demonstrates their collaborative potential in restoration capacity improvement. Additionally, in Fig. 7, we conduct visual comparisons between the models with and without the input-conditional vectors, where the results validate that the latter can better remove the weather. Interestingly, the textual prompt corresponding to the weather scene is beneficial for restoration as it can incorporate explicit scene contexts and semantics, even if it stems from LQ observations. Hence, the model with full C2P achieves the best.
@@
TABLE V: The Model Performance across Three Degradation Scenarios under Varying Vector Lengths.
Vector Number Outdoor-Rain RainDrop Snow100K Average
PSNR SSIM PSNR SSIM PNSR SSIM PSNR SSIM
0 31.95 0.9384 29.87 0.9157 30.67 0.9111 30.83 0.9217
4 31.87 0.9390 30.18 0.9171 30.56 0.9119 30.87 0.9227
8 32.01 0.9405 30.74 0.9206 30.80 0.9142 31.18 0.9251
12 31.72 0.9387 29.13 0.9167 30.27 0.9112 30.37 0.9222
@@
Number of Learnable Vectors. The input-conditional vector in C2P is of the size $N\times C$. To further study the impact of the learnable vectors for adverse weather removal, we initialize the vectors with different numbers (i.e., $N$), where “0” denotes the model without the vector. The results are summarized in Table V. We can see that the performance boosts gradually as the number increases and achieves optimal performance across all degradation types when the vector length is set to 8. This is due to more vectors yielding better capacity for context understanding and weather-related prompt adaption. However, when we increase $N$ from 8 to 12, the model suffers from a performance decrease. This can be attributed to the overlarge vectors introducing unnecessary noisy information that overwhelms other important components. Therefore, we set $N=8$ in our experiments.
@@
In CyclicPrompt, the residual prior is incorporated to enrich visual characteristics for better detail recovery along with EPM. We now present the experimental results to verify its impact. In Table VI, we observe that it contributes to performance gains. When we combine it with the weather-free prompt, one can see that it assists the model well and enhances the performance notably. Moreover, in Fig. 8, we visualize the residual prior map and corresponding feature maps before and after the RPM. There are two observations: (1) the prior (Fig. 8(b)) is weather-invariant and allows expressing the contextualized information of scenes. (2) By modulating the decoder feature (Fig. 8(c)) with the prior, the model can effectively emphasize the textures and structures (Fig. 8(d)), improving the representation capacity for reconstruction.
Impact of Hypaerparameter $\lambda$. We conduct six groups of experiments with the $\lambda$ value ranging from 0.1 to 4, respectively, where the results are illustrated in Table VII. We can see that as $\lambda$ increases within the range of $\leq 1$, the model performance demonstrates observable gains and converges when $\lambda=1$, which indicates that enhancing the impact of the second iteration contributes to improved model effectiveness. Besides, $\lambda>1$ amplifies the impact of the second iteration beyond the first, where the model performance exhibits minimal fluctuations. Hence, we ultimately selected $\lambda=1$ in our experiments.
TABLE VII: The Impact of the Hyperparameter $\lambda$
$\lambda$ 0.1 0.5 1 2 3 4
PSNR 31.06 31.10 31.18 31.17 31.18 31.18
SSIM 0.9243 0.9245 0.9251 0.9253 0.9252 0.9243
Influence of Iteration Number. The proposed CyclicPrompt contains two iterations that perform “Prompt-Restore-Prompt”. To study the influence of the iteration number in UAWR, based on the second iteration, we further extract high-quality image features from the restoration results as restoration priors and perform refinement, thereby constructing model variants with 3 and 4 iterations, respectively. As quantitatively demonstrated in Table VIII, the PSNR metric exhibits insensitivity to the iteration number. While extending the model to 3 iterations yields a marginal SSIM improvement (0.9266 vs. 0.9251), this gain incurs a substantial computational overhead of about 90.5G FLOPs. Furthermore, the model with 4 iterations surprisingly degrades in SSIM with more FLOPs. Consequently, to maintain an optimal balance between efficiency and restoration quality, we set the iteration number to 2 as the final configuration.
TABLE VIII: The Influence of Iteration Number.
Iteration Number PSNR SSIM Flops (G)
2 31.18 0.9251 230.42
3 31.18 0.9266 320.90
4 31.18 0.9248 411.39
IV-E Discussion and Limitations
CyclicPrompt introduces a cyclic prompting framework that integrates C2P and EPM to constitute a cyclic prompting mechanism for UAWR. It involves a two-stage processing (“Prompt-Restore” and “Restore-Prompt”), which suffers from more computational burdens (Table II). Despite its advantages, as shown in Fig. 9, CyclicPrompt struggles to remove weather artifacts in regions where weather patterns and image textures become perceptually indistinguishable. For example, in the 1st row, the rain streaks are more like linear textures on the board, which can cause a misleading understanding of the scene context. Even so, we can see that our method still restores the images better than the recent SotA method GridFormer.
Figure 9: Failure cases of our method under challenging scenarios.